OpenStudy (anonymous):

Write 3^x = 66 as a logarithmic function.

OpenStudy (anonymous):

would it be log(3x)=log(66) log(3) + log(x)=log(66) log(x)=log(66)-log(3) log(x)=log(66/3) log(x)=log(22) x=22

OpenStudy (anonymous):

\[x = \log_{3}66\]

OpenStudy (anonymous):

\[x = \log(66)/\log(3) = 3.81\]

OpenStudy (anonymous):

thank you